OpenStudy (anonymous):

A car has a kinetic energy of 1.9 X 10 cubed joules. If the velocity of the car is decreased by half, what is its kinetic energy?

OpenStudy (noelgreco):

What formula do you use for kinetic energy?

OpenStudy (anonymous):

KE = 1/2 mv squared

OpenStudy (anonymous):

K=(1/2)*m*v^2

OpenStudy (noelgreco):

Right. Now if you cut v in half, you'd have: \[KE _{1} = \frac{ 1 }{ 2 } m (\frac{ 1 }{2 }v)^{2}\] Now, the ony difference is that you're squaring that 1/2 as well as the v. What fraction do you get?

OpenStudy (anonymous):

I dont even know how to get the fraction but the multiple choice answers are A; 2.2 X 10 squared Joules B; 4.7 X 10 squared Joules C; 5.2 X 10 squared Joules D; 5.8 X 10 cubed Joules

OpenStudy (noelgreco):

What is \[\frac{ 1 }{ 2 } \times \frac{ 1 }{ 2 }\] ?

OpenStudy (anonymous):

0.25?

OpenStudy (noelgreco):

good. multiply the original KE by .25 and choose the right answer - it's on there